OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2 A. y - 4 = 1/2(x - 1) B. y + 4 = 2(x - 1) C. y + 4 = 1/2(x - 1) D. y - 4 = 2(x + 1)

OpenStudy (anonymous):

C. y + 4 = 1/2(x - 1)

OpenStudy (anonymous):

the formula is y-y1=m(x-x1)

OpenStudy (anonymous):

thanks!